OpenStudy (anonymous):

ILL GIVE YOU A MEDAL AND FAN How can one fourthx − 3 = one halfx + 8 be set up as a system of equations? How can one fourthx − 3 = one halfx + 8 be set up as a system of equations? How can one fourthx − 3 = one halfx + 8 be set up as a system of equations? How can 1/4x-3=1/2x+8 be set up as a system of equations?

OpenStudy (anonymous):

@texaschic101 sorry it posted the question 30 thousand times lol

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }x-3=\frac{ 1 }{ 2 }+8\]

OpenStudy (anonymous):

@phi @mathlover2014 @mathman101

OpenStudy (texaschic101):

well.....not sure about this...you will need a second opinion I am thinking your set of equations is : y = 1/4x - 3 y = 1/2x + 8 and when solving, you would set it up like this.. 1/4x - 3 = 1/2x + 8

OpenStudy (anonymous):

do you want me to post the answer options?

OpenStudy (texaschic101):

oh...lol...that would help

OpenStudy (anonymous):

alright one sec:)

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

solve for y, then set them equal to each other 4y - x = -12 4y = x - 12 y = 1/4x - 3 2y - x = 16 2y = x + 16 y = 1/2x + 8 therefore, 1/4x - 3 = 1/2x + 8 ......answer is the 2nd one

OpenStudy (anonymous):

thank you so much! i really appreciate it!!!

OpenStudy (texaschic101):

no problem :)